The radius of an atom of rubidium (Rb) is about 2.1 Å. (a) Express this distance in nanometers (nm). Express this distance in picometers (pm). pm (b) How many rubidium (Rb) atoms would have to be lined up to span 1.0 mm? (c) If the atom is assumed to be a sphere, what is the volume in cm3 of a single rubidium (Rb) atom?
Answer #1

Radius of Rubidium (R6) atom = 2.16 la) We know that WE LAO = 10-8 cm =10 lom & luma 107cm = 109m .:1A = 10lom - 10 x 10 9 In(3) 1pm - 10-12m 10 10 x 102mm - 10-2 A la = 100pm 2:14° = ? 2.1A x 100pm - TAO ~ 210pm No. of Rb atoms in Imm: (6) Radius ofc)

Volume of the sphere = 4πr3/3

Given that radius of Rb atom = 2.1×10-8 cm

  4* 22 * (2.1 * 108cm) 7 * 3

= 38.808×10-24 cm​​​​​3

=3.881×10-23 cm​​​​​3